### v1

# Bayesian Intrinsic Groupwise Image Registration: Unsupervised Disentanglement of Anatomy and Geometry

Abstract
This article presents a general Bayesian learning framework for multi-modal groupwise registration on medical images. The method builds on probabilistic modelling of the image generative process, where the underlying common anatomy and geometric variations of the observed images are explicitly disentangled as latent variables. Thus, groupwise registration is achieved through the solution to Bayesian inference. We propose a novel hierarchical variational auto-encoding architecture to realize the inference procedure of the latent variables, where the registration parameters can be calculated in a mathematically interpretable fashion. Remarkably, this new paradigm can learn groupwise registration in an unsupervised closed-loop self-reconstruction process, sparing the burden of designing complex intensity-based similarity measures. The computationally efficient disentangled architecture is also inherently scalable and flexible, allowing for groupwise registration on large-scale image groups with variable sizes. Furthermore, the inferred structural representations from disentanglement learning are capable of capturing the latent anatomy of the observations with visual semantics. Extensive experiments were conducted to validate the proposed framework, including four datasets from cardiac, brain and abdominal medical images. The results have demonstrated the superiority of our method over conventional similarity-based approaches in terms of accuracy, efficiency, scalability and interpretability.
Index Terms:
Bayesian Deep Learning, Disentangled Representation, Groupwise Registration, Multi-Modality, Interpretability
1 Introduction
Groupwise image registration aims to find the hidden spatial correspondence that aligns multiple observations. For medical images, when the observations reflect some common anatomy, their intrinsic structural correspondence, which can be independent of the multi-modal imaging acquisition protocol, is of particular interest. However, conventional methods on multi-modal groupwise registration usually rely on intensity-based similarity measures to iteratively optimize the spatial transformations. For instance, multivariate joint entropy or mutual information were proposed as groupwise similarity measures in [1, 2, 3, 4], followed by the more computationally favourable template-based approaches [5, 6]. Nevertheless, devising proper similarity measures and choosing the correct registration hyperparameters for heterogeneous medical images can be tedious and challenging. The heavy computational burden and the instability in registration accuracy may also prevent real-world applications of the conventional groupwise registration algorithms.
In this work, we seek to establish a new unified and interpretable learning framework for large-scale unsupervised multi-modal groupwise registration. Inspired by the recent progress in disentangled representation learning [7, 8], we propose to disentangle the underlying common anatomy and geometric variations from the observed images, which can be regarded as reverting the data generating process of the observations. To this end, a probabilistic generative model is constructed, where the common anatomy and spatial transformations are disentangled as latent variables. Thus, the problem of groupwise registration (recovery of the intrinsic structural correspondence) devolves to the estimation of the latent posterior distribution, which is then solved via variational inference. Besides, since unsupervised disentangled representation learning is impossible without proper inductive biases [9, 10], we have designed a novel hierarchical variational auto-encoding architecture to realize the inference procedure of the latent variables, where the decoder network complies with the equivariance assumption of the imaging process. Accordingly, groupwise registration is learned in an unsupervised closed-loop self-reconstruction process: i) the encoder extracts the intrinsic structural representations from the multi-modal observations, based on which the common anatomy and the spatial correspondence are estimated; ii) the decoder emulates the equivariant image generative process from the common anatomy, and reconstructs the observations via the inverse spatial transformations.
Our contributions can be summarized as follows:
  • We propose a new learning paradigm for multi-modal groupwise registration based on Bayesian inference and disentangled representation learning. Remarkably, we can achieve registration in an unsupervised closed-loop self-reconstruction process, which spares the burden of designing complex image similarity measures.
[…]
  • Based on the equivariance assumption and the decomposition of the symmetry group actions on the observational space, we prove that under certain conditions, our model can identify the desired registration parameters, which constitutes the theoretical underpinnings of the established framework.
  • Our registration model, while trained using small image groups, can be readily adapted to large-scale and variable-size test groups, significantly enhancing its computational efficiency and applicability.
  • We validated the proposed framework on four publicly available medical image datasets, demonstrating its superiority over similarity-based methods in terms of accuracy, efficiency, scalability and interpretability.
The remainder of the article is organized as follows. Section 2 discusses the related work to this study. Section 3 elucidates the proposed Bayesian inference framework for groupwise registration. Section 4 describes the proposed hierarchical variational auto-encoder architecture for learning and estimation of the latent variables. Section 5 presents the experimental setups and evaluation results of our method on four different public datasets. Section 6 notes implications of the proposed framework and concludes the study. For conciseness, we specify in Table I the major mathematical symbols used in the rest of this paper.
TABLE I: Definition of the major mathematical symbols used in this paper.
Symbol
Description
Symbol
Description
$L$
the number of levels of latent variables
$N$
the number of images in the group
$\bm{Z}$
random variable of the common anatomy
$\bm{U}$
random vector of the image group $\{U_{j}\}_{j=1}^{N}$
$\Omega$
the coordinate space of the common anatomy
$\Omega_{j}$
the $j$-th image coordinate space
$\bm{\phi}$
the set of diffeomorphic maps $\bm{\phi}\triangleq\{\phi_{j}\}_{j=1}^{N}$
$\phi_{j}$
the spatial transformation from $\Omega$ to $\Omega_{j}$
$\bm{z}$
the latent common structural representation
$\bm{u}$
an observed sample of the image group
$\bm{z}^{l}$
the latent common structural representation at the $l$-th level
$\bm{v}^{l}$
the latent stationary velocity fields registering the image group at the $l$-th level
$q^{*}(\bm{z}^{l}\,|\,\bm{u},\bm{v})$
the geometric mean categorical variational distribution of $\bm{z}^{l}$ given $\bm{u}\circ\bm{\phi}\triangleq\{u_{j}\circ\phi_{j}\}_{j=1}^{N}$
$\bm{v}_{j}^{l}$
the stationary velocity field of the $j$-th diffeomorphism at the $l$-th level, $\bm{v}^{l}=\{\bm{v}_{j}^{l}\}_{j=1}^{N}$
$q_{j}^{\diamond}(\bm{z}^{l}\,|\,u_{j},\bm{v}_{j})$
the $j$-th single-view categorical variational distribution of $\bm{z}^{l}$ given $u_{j}\circ\phi_{j}$
$\bm{\mu}_{\bm{v},j}^{l},\bm{\Sigma}_{\bm{v},j}^{l}$
parameters of the Gaussian variational distribution $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l})$
$\widetilde{q}_{j}(\bm{z}^{l}\,|\,u_{j})$
the $j$-th single-view categorical variational distribution of $\bm{z}^{l}$ given $u_{j}$
$\bm{v}_{j}^{+}$
the total velocity field aggregating all $\{\bm{v}_{j}^{l}\}_{l=1}^{L}$ upsampled to the top level
$\bm{\psi}$
parameters of the variational model $\{\bm{\psi}_{j}\}_{j=1}^{N}$
$\bm{\theta}$
parameters of the generative model
2 Related Work
In this section, we review related work to the proposed framework, including similarity-based and deep feature-based approaches to (groupwise) image registration, as well as literature on multi-modal representation learning that facilitates disentanglement of latent variables from multiple modalities. We hope to expose the connections between them that motivate our established framework.
2.1 Groupwise Image Registration
Let $\bm{U}=\{U_{j}\}_{j=1}^{N}$ be the random vector representing the image group and $\bm{u}=\{u_{j}\}_{j=1}^{N}$ an observed sample of $\bm{U}$, with $U_{j}:\mathds{R}^{d}\supset\Omega_{j}\rightarrow\mathds{R}$ the intensity mapping, $\Omega_{j}$ the image domain, and $d$ the dimensionality. In the classical pattern matching and computational anatomy regime [11, 12, 13, 14, 15], the observed homogeneous image group is considered as samples from the orbit of a transformation group $\mathcal{G}$ acting on a deformable template $U_{0}$ (a.k.a. atlas), namely
$\bm{U}\subset\mathcal{G}\cdot U_{0}\triangleq\{U_{0}\circ\phi_{j}^{-1}:\phi_{j}\in\mathcal{G}\},$ (2)
where the transformation $\phi_{j}^{-1}:\Omega_{j}\rightarrow\Omega$ maps spatial locations in the image domain to a common coordinate space $\Omega\subset\mathds{R}^{d}$, and $\mathcal{G}$ is often taken as the group of diffeomorphisms. This concept has motivated the development of population averaging methods that estimate the deformations $\mathcal{G}$ and the template $U_{0}$ simultaneously, based on the observations $\bm{U}$ [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. The resultant transformations encode the structural variability and the template provides a statistical representative of the images. In particular, a priori guess of the template as one of the observed images is proposed in [16, 17, 19], with the template updated by the average deformation in each iteration. Joshi et al. [18] and Bhatia et al. [20] proposed the usage of intensity mean image as the template, which avoid reference selection by estimating transformations from the common space. Allassonnière et al. [21] extended the setup to a Bayesian framework, with the template modelled as a linear combination of continuous kernel functions. Recently, Dalca et al. [25], Ding and Niethammer [26] proposed learning-based template estimation methods, using spatial transformation networks to predict diffeomorphisms parameterized by stationary velocity fields [27, 28]. Nevertheless, groupwise registration based on the deformable template assumption could not readily accommodate the multi-modal nature of medical images.
Multi-modality can introduce additional complexity, as the observed images are not related to an intensity template directly through spatial transformations. In this case, based on the common anatomy assumption of multi-modal medical images, we can instead assume an anatomical or structural representation $\bm{Z}$ from which each observed image is generated through the composition of an imaging functional $f_{j}$ and a spatial transformation $\phi_{j}^{-1}$, i.e.
$U_{j}=f_{j}(\bm{Z})\circ\phi_{j}^{-1}.$ (3)
More importantly, the imaging functional is assumed to be spatially equivariant w.r.t. $\phi_{j}^{-1}$, a condition motivated by the demand to learn equivariant image features [29, 30, 31, 32], which writes
$f_{j}(\bm{Z})\circ\phi_{j}^{-1}=f_{j}(\bm{Z}\circ\phi_{j}^{-1}),\quad\forall\,\phi_{j}\in\mathcal{G}.$ (4)
Inspired by inter-modality pairwise registration based on the joint intensity distribution (JID) or particularly the mutual information (MI) [33, 34, 35, 36, 37], initial attempts to realize multi-modal groupwise registration focused on generalizing this information-theoretic approach directly to high-dimensional cases [1, 2, 38, 3, 4]. Later, concerning the curse of dimensionality in estimating the high-dimensional JIDs, template-based groupwise registration revives through probabilistic modelling of the image generative process in Eq. 3 [39, 40, 41, 5, 42, 6]. These methods assumed that the JID is modelled by the marginalization of the joint distribution between the latent template and the observed images. Thus, the spatial transformations were determined by maximum likelihood estimators (MLEs) while the template was predicted via maximum a posteriori (MAP). We direct interested readers to our previous work [6] for detailed elucidation of this maximum-likelihood perspective and its connection to information-theoretic metrics.
[…]
2.2 Deep Feature-Based Image Registration
[…]
2.3 Multi-Modal Representation Learning
The underlying structural representations that facilitate the estimation of spatial correspondence can be regarded as some latent embedding, which encodes the information correlating the multi-modal observations. This viewpoint, in the context of heterogeneous data, brings the need for multi-modal representation learning which exploits the commonality and complementarity of multiple modalities, and the demand for data alignment which identifies the relation and correspondence between different modalities [55].
Modern probabilistic generative models [56, 57, 58, 59] construct a general framework for representation learning, as the data formation procedure to which they adapt is optimally suited for an unsupervised manner, and they are powerful in striking a balance between fitting and generalizability by taking into account the uncertainty in characterizing the data and latent distributions.
One family of methods within this scope is variational autoencoders (VAEs) [56], where the paradigm is to approximate the intractable latent distribution by a learnable variational posterior during optimizing an evidence lower bound (ELBO) for the log likelihood of observed data. It provides an information-theoretic way to establish the foundational underpinnings of data encoding for downstream tasks [60]. The multi-modal variants of VAEs additionally imbue the capability of retrieving modality-invariant representation by factorizing it into single-view posteriors (a.k.a. experts) inferred from each modality. A typical factorization specifies the combination of the experts explicitly (e.g. mixtures [61], products [62], mixtures of products [63]) or implicitly (e.g. cross-modal generation [64]), compelling the summarization of the common information from any individual modality.
These approaches, nonetheless, all necessitate originally aligned inputs, such as the image and sentence that describe the same object [61], or multi-modal images that are well-registered [65, 66], which limits their applicability to distorted data. Besides, they often adopt a reductionist stance by assuming an overly simplistic prior. (e.g., the standard Gaussians or Laplacians), failing to respect the structure of latent manifold induced by real-world high-dimensional images, which is crucial for mitigating off-track regularization and vanishing latent dimensions [67, 68], as well as for ensuring suitable decomposition of latent encodings [69]. The mixture-based factorizations of the posterior also suffer from inevitable gap between the ELBO and true likelihood, leading to undesirable suboptimal results [70].
In contrast, our framework embraces advancements in tackling all the limitations above: it allows to learn from distorted images in a unified manner, through explicitly endowing the representations of spatial transformations and the common anatomy with a disentangled nature.
3 Bayesian Groupwise Registration
(a) Generative model.
(b) Inference steps #1.
(c) Inference steps #2.
(d) Generation steps.
Figure 1: The proposed hierarchical framework for Bayesian groupwise registration (3-layer example). Random variables are in circles, deterministic variables are in double circles, and observed variables are shaded. Diamonds denotes network feature maps, and squares represent the variational distributions. (a) Probabilistic graphical model of the generative process. (b) Inference steps #1 that predict the hierarchical velocity fields, where we denote $\widetilde{\bm{q}}_{\bm{z}^{l}}\triangleq\{\widetilde{q}_{j}(\bm{z}^{l}\mid u_{j};{\bm{\psi}_{j}})\}_{j=1}^{N}$ and $\bm{q}_{\bm{v}^{l}}\triangleq\{q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})\}_{j=1}^{N}$. (c) Inference steps #2 that predict the common structural representations base on the warped images, where we denote $\bm{q}_{\bm{z}^{l}}^{\diamond}\triangleq\{q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})\}_{j=1}^{N}$ and $\bm{q}_{\bm{z}^{l}}^{*}\triangleq q^{*}(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$. (d) Generation steps that reconstruct the original images. Note that the inference and generation steps form a closed-loop self-reconstruction process.
We propose to formulate the estimation of groupwise registration as a problem of Bayesian inference. Specifically, let $\bm{u}=(u_{j})_{j=1}^{N}$ be a sample of the random vector $\bm{U}$ of the image group. From Eq. 3, each image $u_{j}$ is generated from the latent variable $\bm{z}$ representing the corresponding common anatomy, and the diffeomorphic transformation $\phi_{j}^{-1}$. To ensure its invertibility, we assume that the transformations $\bm{\phi}\triangleq(\phi_{j})_{j=1}^{N}$ is parameterized by stationary velocity fields $\bm{v}=(\bm{v}_{j})_{j=1}^{N}$ [27, 28] such that $\phi_{j}=\exp(\bm{v}_{j})$, and
$\frac{\partial}{\partial t}\phi_{j}(\bm{\omega},t)=\bm{v}_{j}(\phi_{j}(\bm{\omega},t)),\quad\forall\,\bm{\omega}\in\Omega,\ t\in[0,1].$ (5)
The diffeomorphism of the transformations is further imposed by assigning proper priors to the velocity fields.
3.1 Hierarchical Bayesian Inference
In practice, registration is often performed at multiple levels to facilitate convergence of the algorithm [71, 72]. Therefore, we express the latent variables with $L$ hierarchical levels, i.e. $\bm{z}=(\bm{z}^{l})_{l=1}^{L}$ and $\bm{v}_{j}=(\bm{v}_{j}^{l})_{l=1}^{L}$, in which higher levels indicate finer resolutions. Thus, the graphical model of the image generative process can be described as Fig. 0(a). Note that while different levels of the latent variables are assumed to be independent, the inference procedure is performed hierarchically.
Since we are going to parameterize the likelihood $p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})$ by neural networks in the context of Bayesian deep learning, exact inference and parameter estimation becomes intractable [56]. Therefore, we resort to variational inference (VI) to approximate the maximum likelihood. The objective function of VI is the evidence lower bound (ELBO) of the log-likelihood function, i.e.
$\mathcal{L}(\bm{\theta},\bm{\psi}\mid\bm{u})$ $&\mathbb{E}_{q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})}[\log p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})]\\ &-D_{\text{KL}}\big{[}q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})\;\|\;p(\bm{z})p(\bm{v})\big{]},\end{aligned}$ (6)
$\leq\log p(\bm{u};\bm{\theta})\triangleq\ell(\bm{\theta}\mid\bm{u})$
where $q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})$ is the variational distribution, an approximation to the intractable true posterior $p(\bm{z},\bm{v}\mid\bm{u};\bm{\theta})$; $\bm{\psi}$ and $\bm{\theta}$ are the variational and generative parameters respectively. The expectation over the likelihood can be estimated by Monte-Carlo sampling [56].
To simplify the KL divergence term, we first write
$D_{\text{KL}}\big{[}q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})\;\|\;p(\bm{z})p(\bm{v})\big{]}$ (7)
$=\mathbb{E}_{q(\bm{v}\mid\bm{u};\bm{\psi})}\Big{[}D_{\text{KL}}\big{[}q(\bm{z}\mid\bm{u},\bm{v};\bm{\psi})\;\|\;p(\bm{z})\big{]}\Big{]}+D_{\text{KL}}\big{[}q(\bm{v}\mid\bm{u};\bm{\psi})\;\|\;p(\bm{v})\big{]},$
where $q(\bm{v}\mid\bm{u};{\bm{\psi}})=\prod_{l=1}^{L}q(\bm{v}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})$, with $\bm{v}^{l}\triangleq(\bm{v}_{j}^{l})_{j=1}^{N}$, and $\bm{v}^{<l}$ denotes the velocity fields in levels lower than $l$. This factorization over velocity fields can be illustrated by the inference steps #1 in Fig. 0(b). Besides, we assume that the multi-level variational posteriors of the velocity fields factorizes as
$q(\bm{v}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})=\prod_{j=1}^{N}q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}}).$ (8)
Then, the KL w.r.t. the velocity fields can be decomposed as
$D_{\text{KL}}\big{[}q(\bm{v}\mid\bm{u};\bm{\psi})\;\|\;p(\bm{v})\big{]}$ (9)
=\ = $\sum_{j=1}^{N}\sum_{l=1}^{L}\mathbb{E}_{q(\bm{v}_{j}^{<l}\mid\bm{u};{\bm{\psi}})}\left[D_{\text{KL}}\big{[}q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})\;\|\;p(\bm{v}_{j}^{l})\big{]}\right].$
Likewise, the KL w.r.t. the common anatomy can be simplified into
$D_{\text{KL}}\big{[}q(\bm{z}\mid\bm{u},\bm{v};\bm{\psi})\;\|\;p(\bm{z})\big{]}$ (10)
=\ = $\sum_{l=1}^{L}\mathbb{E}_{q(\bm{z}^{<l}\mid\bm{u},\bm{v};{\bm{\psi}})}\left[D_{\text{KL}}\big{[}q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})\;\|\;p(\bm{z}^{l})\big{]}\right],$
where we have assumed that
$q(\bm{z}^{l}\mid\bm{u},\bm{v},\bm{z}^{<l};{\bm{\psi}})=q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}}),$
i.e., the common anatomy at level $l$ can be inferred directly from $\bm{u}$ and $\bm{v}$ without referring to lower-level representations $\bm{z}^{<l}$, which is illustrated by Fig. 0(c).
Figure 2: An example of the geometric and arithmetic mean for the single-view Gaussian posterior distributions.
[…]
Eq. 11 summarizes the decomposition of KL divergence w.r.t. the latent variables. Note that for simplicity we have defined $q(\bm{v}_{j}^{<1}\mid\bm{u};{\bm{\psi}})=q(\bm{z}^{<1}\mid\bm{u},\bm{v};{\bm{\psi}})\triangleq 1$. As will be elucidated in the following subsections, the key idea behind Eq. 11 is: the overall KL divergence is decomposed w.r.t. (i) the common structural representations $\bm{z}$, and (ii) the velocity fields $\bm{v}$. The former is used to measure the intrinsic structural dissimilarity among the observed multi-modal images, while the latter serves as smoothness regularization to enforce diffeomorphism of the transformations.
3.2 Intrinsic Distance over Structural Representations
As mentioned before, the KL divergence w.r.t. the common structural representations is used to measure the dissimilarity of the observed images for registration. Namely, instead of using similarity measures in the image space, we propose to learn an intrinsic distance over the structural representations corresponding to the observed images.
We first extract the single-view structural representation maps corresponding to each individual image, represented by the distributions $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})$, $j=1,\dots,N$, which is predicted from the inference steps #2 (ref. Fig. 0(c)). Note that $\bm{\psi}_{j}$ denotes the modality-specific variational parameters for each image, which are the same for images of the same modality. Then we define the variational posterior of the common anatomy as the geometric mean of these single-view posteriors, i.e.
$q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}}):=q^{*}(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})\propto\left[\prod_{j=1}^{N}q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})\right]^{\nicefrac{{1}}{{N}}},$ (12)
which captures the common information among the multi-modal images. On the other hand, the prior distribution of the common anatomy is given by the arithmetic mean of the single-view posteriors, i.e.
$p(\bm{z}^{l};{\bm{\psi}}):=p^{+}(\bm{z}^{l};{\bm{\psi}})\triangleq\frac{1}{N}\sum_{j=1}^{N}q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}}),$ (13)
which expresses a priori knowledge of the common anatomy by the mixture of experts. This setup is closely related to the variational mixture of posteriors prior (VampPrior) introduced in [67, 68], where the VampPrior with pseudo-inputs is chosen to maximize the ELBO. For example, Fig. 2 illustrates the geometric and arithmetic mean of Gaussian distributions when $N=3$. One can observe that the three individual modes of single-view posteriors collapse to one after computing the geometric mean. Indeed, if the experts are Gaussian, one can prove that the KL divergence from the arithmetic to the geometric mean distribution is minimized when the experts are identical.
On the other hand, applying Jensen’s inequality to the KL divergence yields
$D_{S}$ $\triangleq D_{\text{KL}}\big{[}q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})\;\|\;p^{+}(\bm{z}^{l};{\bm{\psi}})\big{]}$ (14)
$\leq\frac{1}{N}\sum_{j=1}^{N}D_{\text{KL}}\big{[}q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})\;\|\;q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})\big{]}\triangleq\widetilde{D}_{S},$
where $\widetilde{D}_{S}$ is an upper bound of the original (possibly) intractable KL divergence $D_{S}$ involving mixture distributions. Thus, in practice we use this upper bound as a surrogate for minimizing the KL divergence w.r.t. the latent variable $\bm{z}$.
Note that when minimizing $\widetilde{D}_{S}$ w.r.t. the distribution $q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$, we also obtain the geometric mean $q^{*}(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$ in Eq. 12, i.e.
$q^{*}(\bm{z}\mid\bm{u},\bm{v};\bm{\psi})=\operatorname*{arg\,min}_{q}\widetilde{D}_{S}[q],$
where $\widetilde{D}_{S}[q]$ is regarded as a functional of the distribution $q$. This also justifies the usage of geometric mean as the variational posterior of the common anatomy. In fact, as KL divergence is a statistical distance between probability distributions, the $\widetilde{D}_{S}$ in Eq. 14 serves as an intrinsic distance between the structural representations of the common anatomy and each observed image. Therefore, since geometric mean is the distribution that minimizes this intrinsic distance, the optimization of $\widetilde{D}_{S}^{*}\triangleq\widetilde{D}_{S}[q^{*}]$ w.r.t. $q_{j}^{\diamond}$’s will then seek to force the single-view structural representations $q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})$’s (which are conditioned on the transformation variables) to be identical, thus driving the registration process. Appendix A provides more details for the above reasoning.
3.2.1 Structural Representations for Variable Group Sizes
In our previous work [73], we have used Gaussian distributions to parameterize the structural representations. That is, we assume $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\mathcal{N}(\bm{z};\bm{\mu}_{\bm{z},j}^{\diamond},\bm{\Sigma}_{\bm{z},j}^{\diamond})$ where $\bm{\Sigma}_{\bm{z},j}^{\diamond}$ are diagonal, and for notational conciseness we omit the superscript $l$ indicating the level of the latent variables. We can calculate that the geometric mean is also a Gaussian distribution, i.e. $q^{*}(\bm{z}\mid\bm{u},\bm{v};{\bm{\psi}})=\mathcal{N}(\bm{z};\bm{\mu}_{\bm{z}}^{*},\bm{\Sigma}_{\bm{z}}^{*})$, where
$\bm{\Sigma}_{\bm{z}}^{*}=N\left[\sum_{j=1}^{N}\bm{\Sigma}_{\bm{z},j}^{\diamond\,-1}\right]^{-1},\quad\bm{\mu}_{\bm{z}}^{*}=\frac{\bm{\Sigma}_{\bm{z}}^{*}}{N}\sum_{j=1}^{N}\bm{\Sigma}_{\bm{z},j}^{\diamond\,-1}\bm{\mu}_{\bm{z},j}^{\diamond}.$ (15)
Therefore, the intrinsic distance has a closed-form expression, i.e.
$D_{\text{KL}}\big{[}q^{*}(\bm{z}\mid\bm{u},\bm{v};{\bm{\psi}})\;\|\;q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})\big{]}$
=\ = $\frac{1}{2}\Bigg{[}\log\frac{\absolutevalue{\bm{\Sigma}_{\bm{z},j}^{\diamond}}}{\absolutevalue{\bm{\Sigma}_{\bm{z}}^{*}}}+\operatorname{tr}\left(\bm{\Sigma}_{\bm{z},j}^{\diamond\,-1}\bm{\Sigma}_{\bm{z}}^{*}\right)$
$\quad\;+(\bm{\mu}_{\bm{z},j}^{\diamond}-\bm{\mu}_{\bm{z}}^{*})^{\intercal}\bm{\Sigma}_{\bm{z},j}^{\diamond\,-1}(\bm{\mu}_{\bm{z},j}^{\diamond}-\bm{\mu}_{\bm{z}}^{*})\Bigg{]}+\text{const.}$.
where the quadratic term is essentially a Mahalanobis distance between the structural representations.
Here, we extend the framework using categorical distribution to improve interpretability of the structural representations. Particularly, we propose to model them by independent categorical latent variables, i.e., $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\operatorname{Cat}(\bm{z};\bm{\pi}_{j,1}^{\diamond},\dots,\bm{\pi}_{j,K}^{\diamond})$, where $\bm{\pi}_{j,k}^{\diamond}=(\pi_{\bm{\omega},j,k}^{\diamond})_{\bm{\omega}\in\Omega}\in[0,1]^{\absolutevalue{\Omega}}$ and $\sum_{k=1}^{K}\bm{\pi}_{j,k}^{\diamond}=\bm{1}\in\mathds{R}^{\absolutevalue{\Omega}}$. Thus, the geometric mean becomes another categorical distribution
$q^{*}(\bm{z}\mid\bm{u},\bm{v};{\bm{\psi}})$ $=\operatorname{Cat}(\bm{z};\bm{\pi}_{1}^{*},\dots,\bm{\pi}_{K}^{*})=\prod_{\bm{\omega}\in\Omega}\prod_{k=1}^{K}(\pi_{\bm{\omega},j,k}^{*})^{z_{\bm{\omega},k}},$ (16)
with
$\bm{\pi}_{k}^{*}=\frac{\left[\prod_{j=1}^{N}\bm{\pi}_{j,k}^{\diamond}\right]^{\nicefrac{{1}}{{N}}}}{\sum_{k=1}^{K}\left[\prod_{j=1}^{N}\bm{\pi}_{j,k}^{\diamond}\right]^{\nicefrac{{1}}{{N}}}}\in[0,1]^{\absolutevalue{\Omega}},$
and the intrinsic distance takes the form
$D_{\text{KL}}\big{[}q^{*}(\bm{z}\mid\bm{u},\bm{v};{\bm{\psi}})\;\|\;q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})\big{]}=\sum_{k=1}^{K}\bm{\pi}_{k}^{*}\log\frac{\bm{\pi}_{k}^{*}}{\bm{\pi}_{j,k}^{\diamond}}.$ (17)
Note that this intrinsic distance can be applied to image groups with variable group sizes, as the geometric mean can be computed from an arbitrary number of structural representations.
3.2.2 Categorical Reparameterization Using Gumbel-Rao
Unfortunately, the Gumbel-Max trick [74] for sampling discrete random variables is not differentiable w.r.t. its parameterization, making it unsuitable for stochastic gradient estimators. Concurrent works [75, 76] have then introduced the Gumbel-Softmax (GS) distribution as a continuous relaxation of the discrete categorical variable, which admits a biased reparameterization gradient estimator. Here, we use a variant of the GS estimator, namely the Gumbel-Rao (GR) estimator [77], to further reduce the variance in gradient estimation via Rao-Blackwellization.
Specifically, for objective functions like the ELBO, the gradient w.r.t. the distribution parameters of an expectation over a function $f(\bm{z})$ needs to be computed, namely $\nabla_{\bm{\psi}}\mathbb{E}_{q(\bm{z};{\bm{\psi}})}[f(\bm{z})]$, where $f(\bm{z})$ can be the likelihood function of the observed variables. To overcome the challenge of gradient computation with discrete stochasticity, the Straight-Through Gumbel-Softmax (ST-GS) estimator [75, 77] uses a continuous relaxation, giving rise to a biased Monte-Carlo gradient estimator of the form
$\nabla_{\text{ST-GS}}^{\text{MC}}\triangleq\frac{\partial f(\bm{z})}{\partial\bm{z}}\frac{\operatorname{d}\operatorname{softmax}_{\tau}(\bm{g}+\log\bm{\pi}(\bm{\psi}))}{\operatorname{d}\bm{\psi}},$ (18)
where the forward pass in $f(\cdot)$ is computed using the non-relaxed discrete samples. Based on the ST-GS, the GR estimator takes the form as
$\nabla_{\text{GR}}^{\text{MC}}\triangleq\mathbb{E}\left[\nabla_{\text{ST-GS}}^{\text{MC}}\mid\bm{z}\right]\approx\frac{\partial f(\bm{z})}{\partial\bm{z}}\left[\frac{1}{S}\sum_{s=1}^{S}\frac{\operatorname{d}\operatorname{softmax}_{\tau}(\bm{G}^{s}(\bm{\psi}))}{\operatorname{d}\bm{\psi}}\right],$ (19)
where $\bm{G}^{s}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\bm{g}+\log\bm{\pi}\mid\bm{z}$ for $s=1,\dots,S$, and $\bm{g}$ is a random vector with its entries i.i.d. sampled from the Gumbel distribution. More details on the Gumbel-Rao estimator can be found in Appendix B.
3.3 Spatial Regularization for Diffeomorphisms
To impose spatial smoothness of the velocity fields, we define its prior distribution by $p(\bm{v}_{j}^{l})=\mathcal{N}(\bm{v};\bm{0},\bm{\Lambda_{v}}^{-1})$, where the precision matrix $\bm{\Lambda_{v}}$ is given by the scaled Laplacian matrix of a neighbourhood graph on the voxel grid [46], i.e. $\bm{\Lambda_{v}}=\lambda\bm{L_{v}}=\lambda(\bm{D_{v}}-\bm{A_{v}})$, with $\bm{D_{v}}$ and $\bm{A_{v}}$ the degree and adjacency matrices, respectively. On the other hand, the variational posterior of the velocity fields is given by a mean-field Gaussian distribution, i.e. $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})=\mathcal{N}(\bm{v};\bm{\mu}_{\bm{v},j}^{l},\bm{\Sigma}_{\bm{v},j}^{l})$ where $\bm{\Sigma}_{\bm{v},j}^{l}$ is diagonal. Therefore, the KL divergence for the velocity fields has an explicit formula
$D_{\text{KL}}\big{[}q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})\;\|\;p(\bm{v}_{j}^{l})\big{]}$ (20)
=\ = $\frac{1}{2}\left[\operatorname{tr}(\lambda\bm{D_{v}}\bm{\Sigma}_{\bm{v},j}^{l}-\log\bm{\Sigma}_{\bm{v},j}^{l})+\frac{\lambda}{2}\sum_{r}\sum_{q\in\mathcal{N}(r)}(\bm{\mu}_{\bm{v},j}^{l}[r]-\bm{\mu}_{\bm{v},j}^{l}[q])^{2}\right]$
$+\operatorname{const.}$
where $\mathcal{N}(r)$ are the neighbours of voxel $r$. Thus, the quadratic term over $\bm{\mu}$ enforces the velocity fields to be spatially smooth, encouraging diffeomorphic transformations.
4 An Interpretable Registration Model via Bayesian Disentanglement Learning
Figure 3: An overview of the proposed disentangled auto-encoding framework for interpretable groupwise registration.
Figure 4: The network architecture for the proposed Bayesian groupwise registration, composed of the encoders that extract categorical structural representation maps, the registration modules that calculate multi-scale velocity fields, and the decoders that reconstruct the original images based on the common structural representations. Without loss of generality, the illustration is with $L=3$ levels and $N=3$ images to co-register. Note that the inference steps #2 are only performed in the training stage, while in the test stage the encoder is just fed with the original image group to predict groupwise registration. The purple boxes indicate the calculation of related terms in the ELBO.
Most learning-based registration methods in recent literature are based on an end-to-end training pipeline, where a black-box neural network directly predicts the spatial transformation by optimizing image-level similarity measures [44, 45, 46, 78, 79]. However, not only do the end-to-end black-box networks lack interpretability, learning the complex mapping from multi-modal images to their spatial correspondence disregards their underlying structural relationship and may be prone to generalization error [80].
Bayesian deep learning [81], however, could be a remedy. It unifies probabilistic graphical models (PGM) with deep learning, and integrates the inference and perception tasks, enabling them to benefit from each other. Particularly in our context, the perception component includes the encoders that extract single-view posteriors from individual modalities, while the task-specific component comprises the registration modules and reconstruction decoders that learn jointly to infer the spatial correspondence among the input images.
Fig. 3 presents a schematic illustration of the proposed interpretable registration network, specially designed based on the optimization procedure for Bayesian inference outlined in the previous section. The encoders first disentangle the input multi-modal images into structural representation maps with categorical distribution. Then an average representation of the common anatomy is computed within the latent space, so that spatial correspondence between the observed images and the common anatomy are predicted from these structural representations. The disentanglement is learned by means of inversely simulating the image generative process $p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})$ with decoders, i.e. reconstructing the original images based on the common anatomy representation. Fig. 4 depicts the entire network architecture, composed of the encoders, the registration modules and the decoders. The following subsections will detail the construction of these modules.
4.1 Inference of Structural Representations
In the inference steps #1 (ref. Fig. 0(b)), given the original misaligned image group $\bm{u}$, the encoders first estimate the multi-level categorical distributions $\widetilde{q}_{j}(\bm{z}^{l}\mid u_{j};{\bm{\psi}_{j}})$ for each original image $u_{j}$, where $l$ indicates the spatial resolution of the representation. Note that $\widetilde{q}_{j}(\bm{z}^{l}\mid u_{j};{\bm{\psi}_{j}})\triangleq\operatorname{Cat}(\bm{z};\widetilde{\bm{p}}_{j,1}^{l},\dots,\widetilde{\bm{p}}_{j,K}^{l})$ differs from $q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\operatorname{Cat}(\bm{z};\bm{\pi}_{j,1}^{\diamond\,l},\dots,\bm{\pi}_{j,K}^{\diamond\,l})$ regarding the network input. The former takes $\bm{u}=\{u_{j}\}_{j=1}^{N}$ as input, while the latter takes $\bm{u}\circ\bm{\phi}=\{u_{j}\circ\phi_{j}\}_{j=1}^{N}$. The architecture of the encoders is adapted from the Attention U-Net [82], which enhances the vanilla U-Net [83] with additive attention gates. Moreover, domain-specific batch normalization [84] is utilized in the contracting path of the encoders to cancel out modality-specific appearance variations, while all convolutional layers are shared across modalities, embedding multi-modal images into modality-invariant representations.
4.2 Inference of Velocity Fields
Based on the structural representation maps extracted from the original images by the encoders, the registration modules seek to estimate the transformations that spatially register the image group. This is realized in a coarse-to-fine approach, starting from prediction of the lowest-resolution velocity fields, (ref. the red arrows in Fig. 0(b)).
Particularly, denoting $\widetilde{\bm{p}}_{j}^{l}\triangleq\widetilde{q}_{j}(\bm{z}^{l}\mid u_{j};{\bm{\psi}_{j}})$, the registration module at level $l$ takes as input the partially registered single-view structural probability maps $\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l}$ and their geometric mean for the common anatomy $\widetilde{\bm{p}}^{l*}(\bm{\phi}^{<l})$, based on which the velocity field distribution $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})$ is predicted. Note that the spatial transformations $\bm{\phi}^{<l}\triangleq\{\phi_{j}^{<l}\}_{j=1}^{N}$ are integrated by $\phi_{j}^{<l}=\exp(\bm{v}_{j}^{1}+\dots+\bm{v}_{j}^{l-1})$, with $\bm{v}_{j}^{l}$ sampled from $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})=\mathcal{N}(\bm{v};\bm{\mu}_{\bm{v},j}^{l},\bm{\Sigma}_{\bm{v},j}^{l})$. Finally, the total transformations that register the observed image group are given by $\phi_{j}=\exp(\bm{v}_{j}^{+})$ with $\bm{v}_{j}^{+}\triangleq\sum_{l=1}^{L}\bm{v}_{j}^{l}$, where $\bm{v}_{j}^{l}$ is a random sample during training while $\bm{v}_{j}^{l}=\bm{\mu}_{\bm{v},j}^{l}$ for testing. Note that the velocity fields are upsampled to the finest resolution before summation.
As the structural representation maps $\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l}$ encode the structural features of the partially registered input images, we can in fact compute the mean $\bm{\mu}_{\bm{v},j}^{l}$ of the variational posterior $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})$ directly from them based on the Demons algorithm [85, 86, 87, 88]. Particularly, for each grid location $\bm{\omega}$ at level $l$, we define $\bm{\mu}_{\bm{v},j}^{l}$ by
$\bm{\mu}_{\bm{v},j}^{l}(\bm{\omega})\triangleq\alpha\cdot{\widetilde{\bm{\mu}}_{\bm{v},j}^{l}(\bm{\omega})}\Big{/}\max_{\bm{\omega}\in\Omega^{l}}{\norm{\widetilde{\bm{\mu}}_{\bm{v},j}^{l}(\bm{\omega})}_{2}},$
where
$\widetilde{\bm{\mu}}_{\bm{v},j}^{l}(\bm{\omega})\triangleq-\Big{[}\bm{J}_{\bm{\varphi}_{j}^{l}}^{\intercal}(\bm{\omega})\bm{J}_{\bm{\varphi}_{j}^{l}}(\bm{\omega})+\sigma_{\bm{\varphi}_{j}^{l}}^{2}(\bm{\omega})\bm{I}_{d}\Big{]}^{-1}\!\bm{J}_{\bm{\varphi}_{j}^{l}}^{\intercal}(\bm{\omega})\bm{\varphi}_{j,\bm{\omega}}^{l}(\bm{0})$
is the symmetric Demons force, and $\alpha\in(0,\alpha_{0})$ is a learnable parameter controlling the magnitude of the velocity fields. Specifically,
$\bm{\varphi}_{j,\bm{\omega}}^{l}({\bm{u}})\triangleq\widetilde{\bm{p}}^{l*}(\bm{\omega})-\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l}\circ(\operatorname{id}+\bm{u})(\bm{\omega}),$
is the feature difference given a displacement $\bm{u}\in\mathds{R}^{d}$,
$\bm{J}_{\bm{\varphi}_{j}^{l}}(\bm{\omega})\triangleq-\frac{1}{2}\left[\nabla_{\bm{\omega}}^{\intercal}\,\widetilde{\bm{p}}^{l*}(\bm{\omega})+\nabla_{\bm{\omega}}^{\intercal}\,\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l}(\bm{\omega})\right]$
is the symmetrized Jacobian matrix at the origin, and
$\sigma_{\bm{\varphi}_{j}^{l}}^{2}=\frac{1}{\absolutevalue{\Omega^{l}}-1}\sum_{\bm{\omega}\in\Omega^{l}}\norm{\bm{\varphi}_{j,\bm{\omega}}^{l}({\bm{\bm{0}}})-\overline{\bm{\varphi}}_{j}^{l}(\bm{0})}_{2}^{2}$
is the sample variance of the feature difference norm. Detailed derivation can be found in Appendix C, where the Demons algorithm is established from a mathematically interpretable optimization framework. The variance $\bm{\Sigma}_{\bm{v},j}^{l}$ of the variational posterior is predicted by a convolutional block from the concatenation of probability maps $[\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l};\widetilde{\bm{p}}^{l*}]$, whereby the network parameters are shared among different modalities.
In addition, to avoid the degeneracy that deforms the observed images to an arbitrary coordinate space, we constrain the sum of the final velocity fields to be zero by subtracting their average [17], i.e.
$\bm{v}_{j}^{+}\leftarrow\bm{v}_{j}^{+}-\frac{1}{N}\sum_{j=1}^{N}\bm{v}_{j}^{+}.$ (21)
4.3 Bayesian Disentangled Representation Learning
Given the perceptual modules capturing intrinsic structural representations from multi-modal images and their spatial correspondence, we are going to discuss how such mapping functions are learned in the first place. This is achieved by optimizing the network parameters $\bm{\psi}$ and $\bm{\theta}$ in a closed-loop self-reconstruction process that maximizes the ELBO. The single-view structural distributions for estimating the common anatomy are defined by
$q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})\triangleq\widetilde{q}(\bm{z}^{l}\mid u_{j}\circ\phi_{j};{\bm{\psi}_{j}}),$
which is obtained by feeding the transformed images $u_{j}\circ\phi_{j}$ into the encoders in the inference steps #2 (ref. Fig. 0(c)), effectively reverting the image generative process in Eq. 3.
Based on the common structural representations sampled from the geometric mean $q^{*}(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$, the decoders reconstruct the original images in two steps (ref. Fig. 0(d)). The images corresponding to the common anatomy $\widetilde{x}_{j}=f(\bm{z};\bm{\theta}_{j})$ are decoded first. Then, the estimation of the original images $\widehat{\bm{u}}=\{\widehat{x}_{j}\}_{j=1}^{N}$ are obtained by inverse warping, i.e. $\widehat{x}_{j}=\widetilde{x}_{j}\circ\phi_{j}^{-1}$ with $\phi_{j}^{-1}=\exp(-\bm{v}_{j}^{+})$, simulating Eq. 3.
The network architecture of the decoders is given by multi-level convolutional blocks and linear upsampling. At level $l$, the output feature maps from upsampling are multiplied by the sampled categorical variables $\bm{z}^{l}$. Besides, all convolutional layers are assigned with kernel size of 1 isotropically, to suppress spatial correlation and to form spatial equivariance. Thus, the decoded image $\widetilde{x}_{j}$ is in the same spatial orientation as the common anatomy $\bm{z}$, encouraging the registration module to identify the correct velocity fields $\bm{v}_{j}^{+}$ for reconstructing $\widehat{x}_{j}$. This connection between the spatial equivariance and the identifiability of registration is formally established in Appendix E. Furthermore, the batch normalization layers in the decoders are set as the inverse of the counterpart BN functions along the contracting path of the encoders, i.e.
$\bm{c}_{j}^{\text{out}}=\frac{\bm{c}_{j}^{\text{in}}-\bm{\beta}_{m}}{\bm{\gamma}_{m}+{\varepsilon}}\sqrt{\bm{\sigma}_{m}^{2}+\varepsilon}+\bm{\mu}_{m},$
where for each image index $j$ and its corresponding modality $m$, $\bm{c}_{j}^{\text{in}},\bm{c}_{j}^{\text{out}}\in\mathds{R}^{B_{m}\times C\times H\times W\times D}$ are the input and output feature maps, $\bm{\mu}_{m},\bm{\sigma}_{m}\in\mathds{R}^{C}$ are the batch statistics, and $\bm{\beta}_{m},\bm{\gamma}_{m}\in\mathds{R}^{C}$ are the affine parameters. The arithmetic is computed element-wise. In fact, such preference for invertible and spatially decomposable decoders may promote learning of the true disentangled representations [10, 89]. An illustration of the encoder and decoder architecture can be found in Appendix D.
Finally, the expectation $\mathbb{E}_{q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})}[\log p(\bm{u}\mid\bm{z},\bm{v};{\bm{\theta}})]$ is approximated by Monte Carlo with ancestral sampling, where the likelihood is modelled by a Laplace distribution, i.e.,
$p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})$ $=\prod_{j=1}^{N}p(u_{j}\mid\bm{z},\bm{v}_{j};\bm{\theta}_{j})$ (22)
$\propto\prod_{j=1}^{N}\prod_{\bm{\omega}\in\Omega_{j}}\exp\left\{-\frac{\absolutevalue{u_{j}(\bm{\omega})-f(\bm{z};\bm{\theta}_{j})\circ\phi_{j}^{-1}(\bm{\omega})}}{b}\right\},$
where the scale parameter $b$ is set to 1 in all experiments.
The proposed framework also encourages learning of disentangled representations formally defined in [8], which formulates disentangled representations as a reflection on the decomposition of certain symmetry groups acting on the world states (or observations). This explains why we choose to decompose the latent variables of our model into the common anatomy and the corresponding spatial diffeomorphisms: because they reflect two separate symmetry structures of the observed images, i.e., the ontological transformation changing the underlying anatomy, and the diffeomorphic transformation characterizing the geometric variation. More thorough discussion on disentangled representations and its identifiability can be found in Appendix E.
4.4 Learning, Inference and Scalability
Learning of our model is performed in an end-to-end fashion. All model parameters, including the auto-encoders and the registration modules, are trained simultaneously by stochastic gradient ascent to maximize the ELBO. At the inference stage, the transformations that register the image group is given by $\widehat{\bm{\phi}}=\{\widehat{\phi}_{j}\}_{j=1}^{N}$, where $\widehat{\phi}_{j}=\exp(\widehat{\bm{\mu}}_{j}^{+})$ and $\widehat{\bm{\mu}}_{j}^{+}$ is aggregated from the means of the velocity distributions at all levels predicted by the registration modules.
Remarkably, the proposed disentanglement of anatomy and geometry endows our model with powerful scalability. Since geometric and arithmetic averaging can be performed on any number (greater than 1) of representations, our model is capable of processing image groups of arbitrary sizes during training and test. Let the number of images in an input group be $N^{\prime}=\sum_{m=1}^{M}{N_{m}}$, where $N_{m}$ is the number of images of modality $m$. Each image of modality $m$ is processed by the encoder for the corresponding modality to extract structural representations, and then the geometric and arithmetic means are calculated on the representations of all images to estimate the KL divergence and predict deformations for each image. For decoding, the reconstruction from the decoder of modality $m$ is inversely warped for $N_{m}$ times by the deformations for the $N_{m}$ images of modality $m$ to reconstruct original images. Thus, our model allows $N_{m}$’s to be different or even 0, and only requires $N^{\prime}\geq 2$ (for calculating averages) and at least two modalities are available in each group, i.e., $\sum_{m}\text{1}(N_{m}>0)\geq 2$ (for effectively learning modality-invariant representations).
Therefore, our model can handle groups with missing modalities or variable sizes, which are common occurrences in real-world scenarios. Particularly, the scalability of our model is of great value in two aspects:
  • During training, we can drop several modalities/images in each training group, thereby reducing the computational cost and/or allowing for learning from less data effectively. Typically, there are two experimental settings: 1) complete learning, the standard setting, where each group consists of all modalities, with one image per modality, 2) partial learning, where each group only contains two modalities, with one image per modality. Experiments in both settings are discussed in Section 5.
  • During evaluation, our model can register groups with arbitrary sizes, enabling large-scale and variable-size multi-modal groupwise registration, as explored in Section 5.6. In contrast, conventional iterative methods optimize similarity metrics with computational complexities that escalate rapidly with group size. Deep learning baselines, on the other hand, often have fixed channel numbers to take as input the concatenation of images [90], necessitating groups to be of uniform size. These limitations significantly restrict their applicability.
5 Experiments and Results
We evaluated our framework on a total of four publicly available datasets, including BraTS-2021 [91, 92, 93], MS-CMRSeg [94], Learn2Reg Abdominal MR-CT [95, 96, 97], and the OASIS dataset [98]. The experiments were implemented in PyTorch [99] and conducted on $\text{NVIDIA}^{\circledR}$ $\text{RTX}^{\textnormal{TM}}$ 3090 GPUs. The experimental materials, baselines, evaluation metrics and results are detailed as follows.
5.1 Materials
To show that our method applies to various groupwise registration tasks, especially in multi-modal scenarios, we validated it on four different datasets:
  1. 1.
    MS-CMRSeg. This dataset provides multi-sequence cardiac MR images from 45 patients, including LGE (Late Gadolinium Enhanced), bSSFP (balanced-Steady State Free Precession), and T2-weighted scans. The images were preprocessed by affine co-registration, slice selection, resampling to $1\times 1\ \text{mm}$ and ROI extraction, giving 39, 15 and 44 slices for training, validation and test, respectively. Additional misalignment simulated by random FFDs using isotropic control point spacings of $5/10/20/40\ \text{mm}$ were made to further demonstrate registration. The ROI regions were obtained by dilating the foreground mask using a circular filter of 15-pixel radius. Evaluation was conducted on the warped manual segmentations of the myocardium, left and right ventricle. The major challenge of this dataset comes from the complex intensity patterns of the cardiac region, which can misguide the registration.
  2. 2.
[…]
  3. 3.
    Learn2Reg Abdominal MR-CT. This dataset collects 3D T1-weighted MR and CT abdominal images. The data were resampled to $3\times 3\times 3\ \text{mm}$ and cropped to size of $112\times 96\times 112$. Training was performed on the unpaired 40 MR and 50 CT images [97, 96], while 8 MR-CT pairs (16 images) were used for test [100]. The masks provided by the dataset were used to confine the information used for registration. Evaluation was performed on the warped manual segmentations of the liver, spleen, left and right kidneys. The major challenges of this dataset are the large deformation and missing correspondences between the two modalities, as well as the distribution shift between training and test datasets.
  4. 4.
    OASIS. The OASIS-1 dataset contains 414 T1-weighted 3D MR scans from young, middle-aged, non-demented and demented older adults [98]. The images were skull-stripped, bias-corrected, and registered into an affinely-aligned, common template space with FreeSurfer [101, 102]. Evaluation was conducted on the warped manual segmentations of the cortex, subcortical grey matter, white matter, and CSF. The volumes were resampled $2\times 2\times 2\ \text{mm}$ and cropped to size of $80\times 96\times 96$. Besides, 287/40/87 images were randomly selected for training/validation/test. The major challenge of this task comes from the large inter-subject variability of the brain structures.
5.2 Compared Methods
We compared two types of baselines with our models in the experiments. The first type features state-of-the-art iterative methods for multi-modal groupwise registration, including the accumulated pairwise estimates (APE) [4], conditional template entropy (CTE) [5], and $\mathcal{X}$-CoReg [6]. These methods do not need training, but they can be prone to image artifacts, and can hardly register large image groups due to excessive computational burdens. The second type of baselines extends the first type, aiming to learn the groupwise spatial correspondence via a neural network using the aforementioned similarity measures. To promote fair comparison, we used the Attention U-Net [82] for the network backbone, the same architecture as the encoder in our proposed framework. However, note that these learning-based baselines can only be trained to register images of a fixed group size, and therefore cannot be applied to data of variable group sizes in the test stage.
For the proposed model, we implemented two variants of the registration module and adopted two training strategies. Particularly, the registration modules can be a convolutional network as in our previous work [73], or can be replaced by the Demons force proposed in this work. In addition, we compare the partial and complete learning strategies for our model. The partial learning strategy trains the network with only random image pairs, while the trained model can be used to register images of complete group sizes. Thus, the four variants of our model are:
  • Ours-PN: with partial learning and convolutional network based registration modules.
  • Ours-CN: with complete learning and convolutional network based registration modules.
  • Ours-PD: with partial learning and Demons based registration modules.
  • Ours-CD: with complete learning and Demons based registration modules.
Here, “C” and “P” signify Complete and Partial learning strategies, respectively; “N” and “D” indicate Network or Demons based registration modules, respectively.
5.3 Implementation Details
During preprocessing, the intensity range of each image was linearly normalized to $[0,1]$. For the cardiac and abdominal images, they were further multiplied by the ROI mask to confine the information used by the network to predict registration. The encoder has an Attention U-Net architecture with $L=5$ levels, and the $l$-th level consists of two Conv-BN-LeakyReLU blocks with kernel size 3 and $16\times 2^{L-l}$ channels. The output of each convolutional block in the upsampling path is passed to a Conv layer to produce logits of the categorical structural distribution, with $8\times 2^{L-l}$ channels. Samples from the geometric mean structural distribution are passed to the decoder to reconstruct the original images. The decoder is composed of Conv-BN-LeakyReLU-Upsample blocks with kernel size 1 and $8\times 2^{L-l}$ channels.
To facilitate learning, different weights were used for the reconstruction loss, intrinsic structural distance, and registration regularization, corresponding to a likelihood-tempered ELBO [103]. Table II presents the balancing weights of the loss terms for different training datasets. We also added a local normalized cross-correlation (LNCC) term [104] to improve reconstruction on brain images, with its weight marked by values in the parentheses. The hyperparameter for spatial regularization was set to $\lambda=10$ as in [46], and the maximum magnitude of the velocity computed by the Demons algorithm was set to $\alpha_{0}^{l}=10\times 2^{l-L}$ at each level. The number of samples for the GR estimator was set to $S=3$ in all experiments. The network optimization was performed by stochastic gradient descent using the Adam optimizer [105]. Besides, we used learning rate of $1\times 10^{-3}$, and batch size of 20 or 2 for the MS-CMRSeg or the other datasets, respectively.
TABLE II: This table presents the respective weights of the reconstruction loss, intrinsic structural distance and registration regularization for different experimental setups.
Method
BraTS
MS-CMR
Learn2Reg
OASIS
Ours-PN
120 (60), 75, 15
120, 160, 10
N/A
N/A
Ours-CN
120 (60), 75, 20
120, 160, 10
120, 80, 4
120 (20), 40, 4
Ours-PD
120 (60), 120, 8
120, 160, 10
N/A
N/A
Ours-CD
120 (60), 60, 8
120, 160, 10
120, 80, 1
120 (20), 20, 0.1
5.4 Evaluation Metrics
For each test image group in an experiment, groupwise semantic evaluation metrics can be constructed by averaging the pairwise version over all possible pairs of the propagated segmentation masks. In other words, a metric on the $n$-th group $\{u_{i}\}_{i=1}^{N_{n}}$ is calculated as
$\operatorname{Eva}_{n}\triangleq\frac{1}{\binom{N_{n}}{2}}\sum_{\begin{subarray}{c}1\leq i,j\leq N_{n}\\ i\neq j\end{subarray}}\operatorname{Eva}(y_{i}\circ\widehat{\phi}_{i},y_{j}\circ\widehat{\phi}_{j}),$
where $\operatorname{Eva}$ is the Dice similarity coefficient (DSC) or average symmetric surface distance (ASSD), $y_{i},y_{j}$ are distorted masks associated with their respective images $u_{i},u_{j}$, and $\widehat{{\phi}}_{i},\widehat{{\phi}}_{j}$ are the corresponding predicted transformations. The masks are used for model evaluation only.
For experiments on the BraTS-2021 dataset, since the ground-truth spatial correspondence is available, the groupwise warping index (gWI) can be implemented as an additional metric, which measures the root mean squared error for unbiased deformation recovery based on the ground-truth and predicted deformation fields [6], i.e.
$\operatorname{gWI}_{n}\triangleq\frac{1}{N_{n}}\sum_{i=1}^{N_{n}}\sqrt{\frac{1}{\left|\widehat{\Omega}_{i}^{f}\right|}\sum_{\bm{\omega}\in\widehat{\Omega}_{j}^{f}}\left\|\bar{r}_{j}(\bm{\omega})\right\|_{2}^{2}},$ (23)
where $\widehat{\Omega}_{j}^{f}\triangleq\left\{\bm{\omega}\in\Omega\mid\phi_{j}^{\dagger}\circ\widehat{\phi}_{j}(\bm{\omega})\in F\right\}$ with $F$ the foreground region of the initial phantom before distortion and $\phi_{j}^{\dagger}$ the ground-truth deformation for $u_{j}$, and
$\bar{r}_{j}(\bm{\omega})\triangleq r_{j}(\bm{\omega})-\frac{1}{N_{n}}\sum_{j^{\prime}=1}^{N_{n}}r_{j^{\prime}}(\bm{\omega}),\quad r_{j}(\bm{\omega})\triangleq\phi_{j}^{\dagger}\circ\widehat{\phi}_{j}(\bm{\omega})-\bm{\omega}.$ (24)
The gWI provides a fine-grained measurement for voxel-wise co-registration accuracies, whereas the semantic evaluation metrics attend to high-level structural concurrence of distinct tissues. Besides, since all data have isotropic physical spacing, we measured the ASSD and gWI in voxel units for convenience.
5.5 Multi-Modal & Intersubject Groupwise Registration
5.5.1 Experimental Design
This experiment aims to showcase our model’s performance under various data conditions using the four datasets. These conditions include multi-modal data (all datasets except for OASIS), heterogeneous data such as BraTS-2021 of brain tumors and MS-CMRSeg of myocardium infarction, and intersubject groupwise registration (OASIS). Additionally, we aimed to demonstrate the efficiency of the partial learning strategy for our model. To this end, we examined four variants of our model, namely Ours-CN, Ours-CD, Ours-PN, and Ours-PD.
In this experiment, the test groups of each dataset maintained their original sizes, which was equal to the training group sizes used for both the baselines and complete learning of our model. However, in the case of partial learning for our model, the training group size was consistently set to 2, achieved by randomly selecting images before the experiment, rather than in each training iteration. This resulted in a reduced number of training images, posing additional challenges for model generalizability. Note that our model does not have a partial learning version for the Learn2Reg dataset (as each original group is an MR-CT pair) and the OASIS dataset (as it is mono-modal).
5.5.2 Quantitative Results of Registration Accuracy
TABLE III: Evaluation metrics of the groupwise registration results on the MS-CMRSeg, BraTS-2021, Learn2Reg, and OASIS datasets. The top and second-best results for each dataset are highlighted in bold and underline, respectively. The ASSDs were measured in voxel units. The parameter counts are expressed in millions. The $p$-values were computed using the gWIs or DSCs (for the BraTS-2021 and other datasets, respectively) between the method Ours-CD and the others with a two-sided paired $t$-test. |$\absolutevalue{\det J_{\phi}\leq 0}$ | represents the proportion (in %) of voxels with negative Jacobian determinants in the predicted displacements, where the values were first calculated for the foreground region of each registered image and then averaged over all images among all test groups.
Figure 25: Categorical structural representations extracted by the proposed models on an image group from the BraTS dataset.
Figure 26: Categorical structural representations extracted by the proposed models on an image group from the MS-CMRSeg dataset.
Figure 27: Categorical structural representations extracted by the proposed models on an image group from the Learn2Reg Abdominal MR-CT dataset.
Figure 28: Categorical structural representations extracted by the proposed models on an image group from the OASIS dataset.
Method
MS-CMRSeg BraTS-2021
DSC $\uparrow$
ASSD $\downarrow$
#Params
$p$-value
gWI $\downarrow$
DSC $\uparrow$
ASSD $\downarrow$
#Params
$p$-value
None
.$722\pm.104$
$3.86\pm 1.43$
N/A
$<10^{-10}$
$1.430\pm 0.644$
.$610\pm.150$
$0.93\pm 0.42$
N/A
$<10^{-10}$
[…]
APE [4]
.$746\pm.101$
$3.48\pm 1.36$
0.154
$<10^{-10}$
$0.629\pm 0.141$
.$\underline{.707\pm.069}$
$\bm{0.62\pm 0.12}$
7.37
$1.3\times 10^{-1}$
CTE [5]
.$766\pm.100$
$3.15\pm 1.32$
0.154
$<10^{-10}$
$1.223\pm 0.255$
.$500\pm.102$
$1.27\pm 0.45$
7.37
$<10^{-10}$
$\mathcal{X}$-CoReg [6]
.$757\pm.107$
$3.31\pm 1.40$
0.154
$<10^{-10}$
$0.728\pm 0.196$
.$698\pm.086$
$0.65\pm 0.17$
7.37
$<10^{-10}$
[…]
APE-Att
.$799\pm.061$
$2.78\pm 0.72$
8.04
$<10^{-10}$
$0.757\pm 0.153$
.$690\pm.078$
$0.67\pm 0.14$
22.95
$<10^{-10}$
CTE-Att
.$820\pm.066$
$2.45\pm 0.74$
8.04
$<10^{-10}$
$0.916\pm 0.210$
.$661\pm.094$
$0.75\pm 0.20$
22.95
$<10^{-10}$
[…]
Ours-PN
.$803\pm.062$
$2.68\pm 0.70$
5.22
$<10^{-10}$
$\underline{0.608\pm 0.115}$
.$670\pm.071$
$0.69\pm 0.13$
15.10
$<10^{-10}$
Ours-CN
.$\bm{.842\pm.051}$
$\bm{2.17\pm 0.58}$
5.22
$<10^{-10}$
$\bm{0.538\pm 0.108}$
.$\bm{.709\pm.063}$
$\underline{0.62\pm 0.15}$
15.10
$<10^{-10}$
Ours-PD
.$799\pm.060$
$2.74\pm 0.65$
2.85
$<10^{-10}$
$0.720\pm 0.151$
.$670\pm.072$
$0.70\pm 0.14$
15.06
$<10^{-10}$
Ours-CD
.$\underline{.836\pm.043}$
$\underline{2.21\pm 0.47}$
2.85
N/A
$0.663\pm 0.129$
.$669\pm.074$
$0.70\pm 0.15$
15.06
N/A
Method
Learn2Reg OASIS
DSC $\uparrow$
ASSD $\downarrow$
#Params
$p$-value
|$\absolutevalue{\det J_{\phi}\leq 0}$ | (%) $\downarrow$
DSC $\uparrow$
ASSD $\downarrow$
#Params
$p$-value
None
.$415\pm.160$
$5.27\pm 3.01$
N/A
$2.4\times 10^{-4}$
N/A
.$614\pm.027$
$0.96\pm 0.10$
N/A
$<10^{-10}$
[…]
APE [4]
.$554\pm.384$
$6.14\pm 7.66$
10.52
$9.7\times 10^{-2}$
.$5297\pm.0814$
.777±.030plus-or-minus.777.030.$7$030
$0.59\pm 0.09$
8.85
$7.9\times 10^{-8}$
CTE [5]
.$514\pm.373$
$6.85\pm 7.89$
10.52
$5.1\times 10^{-2}$
.$2291\pm.0447$
.$746\pm.031$
$0.62\pm 0.09$
8.85
$<10^{-10}$
$\mathcal{X}$-CoReg [6]
.$664\pm.361$
$6.86\pm 14.1$
10.52
$3.0\times 10^{-1}$
.$1330\pm.0404$
.777±.017plus-or-minus.777.017.$7$017
$0.54\pm 0.05$
8.85
$<10^{-10}$
[…]
APE-Att
.$687\pm.092$
$2.09\pm 0.57$
22.95
$2.9\times 10^{-3}$
.$0479\pm.0130$
.777±.018plus-or-minus.777.018.$7$018
$0.57\pm 0.05$
22.95
$<10^{-10}$
CTE-Att
.$679\pm.069$
$2.17\pm 0.56$
22.95
$1.4\times 10^{-2}$
.$0552\pm.0154$
.$773\pm.039$
$0.59\pm 0.13$
22.95
$4.2\times 10^{-6}$
[…]
Ours-CN
.$\bm{.803\pm.084}$
$\bm{1.32\pm 0.57}$
15.10
$9.3\times 10^{-2}$
.$1746\pm.0305$
.$\bm{.803\pm.007}$
$\bm{0.49\pm 0.02}$
13.26
$<10^{-10}$
Ours-CD
.$\underline{.793\pm.086}$
$\underline{1.39\pm 0.60}$
15.06
N/A
.$\bm{.0066\pm.0029}$
.$\underline{.791\pm.008}$
$\underline{0.51\pm 0.02}$
13.22
N/A
Section 5.5.2 summarizes the means and standard deviations of the evaluation metrics (DSC, ASSD, and gWI) on the four datasets before and after groupwise registration by different methods. The effectiveness of the iterative and deep learning baselines varies across datasets. For example, all iterative approaches show inferior performance on MS-CMRSeg and Learn2Reg compared to the deep learning baselines. Particularly on Learn2Reg, the evaluation reveals that APE and CTE yield marginal improvements in mean DSC, and all iterative methods even result in worse mean ASSDs after registration. Besides, the deep learning baselines performed worse than APE and $\mathcal{X}$-CoReg on BraTS-2021, and all the baselines (except for CTE) exhibit similar performance on OASIS. These results showcase the instability of both types of baselines under complex image conditions. Besides, their performance was achieved at the expense of either time-consuming iterative optimization for each test group, or a substantial increase in parameter count for deep learning baselines.
In contrast, our model with complete learning attained notable improvements in registration accuracy, consistently outperforming all other methods across all datasets. For example, the two variants of our model demonstrate an over 15% enhancement in DSC and an over 33% reduction in ASSD on Learn2Reg, compared with the best baseline. On BraTS-2021, our model also surpasses all baselines, especially in gWI. It is important to highlight that gWI assesses the registration accuracy for all spatial locations within the images, providing a comprehensive evaluation, while the DSC and ASSD are only based on tumor region. Therefore, the results showcase the superiority and versatility of our model in aligning various brain structures with normal and pathological tissues and diverse imaging patterns.
Remarkably, the proposed method also exhibits a significant reduction in parameters compared to the deep learning baselines. For instance, when deploying a CNN and the Demons algorithm in the registration module on MS-CMRSeg, the reductions are approximately 35% and 65%, respectively, and both versions of our model feature a nearly 42% reduction on OASIS.
In addition, the partial learning strategy allows our model to reduce the training image group size to 2, while maintaining performance equal to or surpassing that of the baselines. This reduction results in a 33%/50%/50% (for the MS-CMRSeg/BraTS-2021/OASIS datasets) decrease in training image resources and approximately the same ratio of reduction in memory footprint during training. The results illustrate that our model, trained with partial learning and reduced computational costs on a limited number of images, achieves a level of generalizability comparable to the baselines that leverage the entire dataset. This showcases the potential of our approach in scenarios involving image groups with absent modalities and constrained resources.
Moreover, examining the proportions of voxels with negative Jacobian determinants in the displacements predicted by various methods on the OASIS dataset reveals that all methods exhibit decent smoothness. Notably, Ours-CD achieved a proportion significantly lower than other methods, highlighting the advantage of the proposed registration modules with the Demons algorithm in predicting diffeomorphic transformations. Readers are referred to Appendix F for the proportion values on all the four datasets.
[Uncaptioned image] Xinzhe Luo received the BS degree in mathematics from Fudan University in 2019. He is currently working toward the PhD degree in statistics at the School of Data Science, Fudan University, advised by Prof. Xiahai Zhuang. His research interests include machine learning, medical image analysis, and interpretable deep learning.
[Uncaptioned image] Xin Wang received the BS degree in electrical engineering from Fudan University in 2020. He is currently pursuing a PhD degree in electrical and computer engineering at the University of Washington, supervised by Prof. Linda Shapiro and Prof. Chun Yuan. His research is centered around medical image analysis and machine learning, with a focus on statistical models for multi-modal image fusion.
[Uncaptioned image] Linda Shapiro is the Boeing Endowed Professor in Computer Science and Engineering, Professor of Electrical and Computer Engineering, and Adjunct Professor of Medical Education and Biomedical Informatics at the University of Washington. Her research interests include computer vision, image database systems, pattern recognition, and medical imaging. She is a fellow of the IEEE and the IAPR, and a past chair of the IEEE Computer Society Technical Committee on Pattern Analysis and Machine Intelligence.
[Uncaptioned image] Chun Yuan is a professor in the Department of Radiology and Imaging Sciences at the University of Utah, and a professor emeritus in the Department of Radiology at the University of Washington. He is a fellow of the ISMRM, AHA, and AIMBE. He has pioneered multiple high-resolution MRI techniques to detect vulnerable atherosclerotic plaques and led numerous MRI studies examining carotid atherosclerosis. He is a member of the editorial board for the following Journals: JACC CV imaging, Journal of Cardiovascular MR, and the Journal of Geriatric Cardiology.
[Uncaptioned image] Jianfeng Feng is the chair professor of Shanghai Center for Mathematical Sciences, and the Dean of Institute of Science and Technology for Brain-Inspired Intelligence and School of Data Science in Fudan University. He has been developing new mathematical, statistical and computational theories and methods to meet the challenges raised in neuroscience and mental health researches. Recently, his research interests are mainly in big data analysis and mining for neuroscience and brain diseases. He was awarded the Royal Society Wolfson Research Merit Award in 2011, as a scientist ‘being of great achievements or potentials’, and the Humboldt Research Award in 2023, for his significant contribution in psychiatry research and pioneering work in Digital Twin Brain.
[Uncaptioned image] Xiahai Zhuang is a professor at the School of Data Science, Fudan University. He graduated from Department of Computer Science, Tianjin University, received Master degree from Shanghai Jiao Tong University and Doctorate degree from University College London. His research interests include interpretable AI, medical image analysis and computer vision. His work won the Elsevier-MedIA 1st Prize and Medical Image Analysis MICCAI Best Paper Award 2023.
Figure 25: Categorical structural representations extracted by the proposed models on an image group from the BraTS dataset.
Figure 26: Categorical structural representations extracted by the proposed models on an image group from the MS-CMRSeg dataset.
Figure 27: Categorical structural representations extracted by the proposed models on an image group from the Learn2Reg Abdominal MR-CT dataset.
Figure 28: Categorical structural representations extracted by the proposed models on an image group from the OASIS dataset.